OpenStudy (anonymous):

At noon the outside temperature was –5°F. By 4 p.m. the same day, the temperature had dropped 13°F. What was the outside temperature at 4 p.m.? At 4 p.m. the outside temperature was °F. 7

OpenStudy (anonymous):

-5-13=?

OpenStudy (anonymous):

i got 8

OpenStudy (anonymous):

no when same sign ,add -5-13=-18 when different sign subtract and keep the sign of larger term. -5+13=+8 +5-13=-8 -5-13=-18 5+13=18

OpenStudy (anonymous):

what>

OpenStudy (anonymous):

terms with same sign are added terms with different sign are subtracted. for this question ,it is -18